Daneka and Kisha both opened a savings account on the same day. Faneca deposited $864 into her account, and Kisha deposited $253 into her account. After 6 months, daneka's account increased to $1,106 and Kisha's account increased to $329. What was the percent of increase for each saving account over the 6 past months? Who made a greater percent of change?
Answer by stanbon(75887)

Daneka and Kisha both opened a savings account on the same day. Faneca deposited $864 into her account, and Kisha deposited $253 into her account. After 6 months, daneka's account increased to $1,106 and Kisha's account increased to $329. What was the percent of increase for each saving account over the 6 past months? Who made a greater percent of change?
-----------
Daneka DATA:
1106-864 = 242
242 = 864*r*(1/2)
864*r = 484
rate = 0.56
------------------
Kisha DATA:
329-253 = 76
76 = 253*r*(1/2)
152 = 253*r
r = 0.60
====
Ans: Kisha
---------
Cheers,
Stan H.
----------